E. Graphical Method; 1. 2x-3y=7;3x+y=5 2. 2x-5y=9;3x+4y=25


1
Expert's answer
2022-03-16T17:49:46-0400

"1.\\
\left\{ \begin{matrix}2x-3y=7\\
3x+y=5\end{matrix}\right.\\
2x-3y=7\\
\begin{matrix}
 x &2 &5 \\
 y & -1&1
\end{matrix}\\
3x+y=5\\
\begin{matrix}
 x &2 &1 \\
 y & -1&2
\end{matrix}\\"


"(2,-1)"

"2.\\
\left\{ \begin{matrix}2x-5y=9\\
3x+4y=25\end{matrix}\right.\\
2x-5y=9\\
\begin{matrix}
 x &2 &-3 \\
 y & -1&-3
\end{matrix}\\
3x+4y=25\\
\begin{matrix}
 x &7 &3 \\
 y & 1&4
\end{matrix}\\"




"(7,1)"